Zaheeda scored 9 goals in 5 netball matches At this rate about how many goals will she score in 10 games

1
Expert's answer
2022-01-31T18:04:03-0500

After 5 games the team scored "1.8*5=9" goals.

After 6 games the team scored "9+3=12" goals.

The mean after 6 games is "\frac{12}{6}=2" goals.